OpenStudy (vera_ewing):

Pre-calculus word problem 2/3

OpenStudy (vera_ewing):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

oh pellet the answer in last part was years not seconds dear

OpenStudy (xapproachesinfinity):

i was wondering carbon takes long time to decay

OpenStudy (xapproachesinfinity):

naturally it takes a substance years to decay sorry lol

OpenStudy (vera_ewing):

Oh hahahaha so 13304.65 years?

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (vera_ewing):

Ok, and how would I do this biology one?

OpenStudy (xapproachesinfinity):

let's do this one

OpenStudy (xapproachesinfinity):

rt=ln2 we are looking for t here so t=ln(2)/r

OpenStudy (xapproachesinfinity):

we are given the r=5% per hour

OpenStudy (xapproachesinfinity):

so r=0.05

OpenStudy (xapproachesinfinity):

t=ln2/0.05

OpenStudy (vera_ewing):

So what exactly is the equation?

OpenStudy (vera_ewing):

And what am I solving for?

OpenStudy (xapproachesinfinity):

you are solving for t which is times it take for the bacteria to double (population) equation was given to you rt=ln2 solve for t=ln2/r r is given r=growth rate they specifically gave us r=5%=0.05

OpenStudy (xapproachesinfinity):

so t=ln2/0.05 in hours

OpenStudy (xapproachesinfinity):

i got t=13.86 hours nearly 14 hours

OpenStudy (vera_ewing):

So this is it?

OpenStudy (xapproachesinfinity):

yeah

OpenStudy (vera_ewing):

So it takes 13.87 hours?

OpenStudy (xapproachesinfinity):

so to answer this we say it take nearly 14 hours for a colony of bacteria to double

OpenStudy (xapproachesinfinity):

or just write the number you found 13.86 hours

OpenStudy (vera_ewing):

Got it, thanks! Can you help me with one more? :)

OpenStudy (xapproachesinfinity):

yeh no problem so far easy word problems hahhaha

OpenStudy (vera_ewing):

lol yeah it's just pre-calculus

OpenStudy (xapproachesinfinity):

i makes sense the the bacteria doubles in that much of hours have you taken any bio class?

OpenStudy (vera_ewing):

Taking AP Bio this year

OpenStudy (xapproachesinfinity):

oh good! i have yet to take bio

OpenStudy (vera_ewing):

It's an interesting course, I really like it! :)

OpenStudy (xapproachesinfinity):

yeah science are all interesting stuff once you know it more

OpenStudy (xapproachesinfinity):

did you post?